# Quasi-two-dimensional soliton in a self-repulsive spin-orbit-coupled dipolar binary condensate

### v1

@@ III.1 Quasi-2D SO-coupled binary self-repulsive nondipolar BEC soliton
Figure 1: Contour plot of density $n_{j}$ of a quasi-2D nondipolar ($d=0$) self-repulsive multi-ring $(\mp 1,0)$-type binary Rashba or Dresselhaus SO-coupled BEC soliton for components (a) $j=1$ ($n_{1}$), (b) $j=2$ ($n_{2}$) and (c) total density ($n$); the same of a stripe soliton for components (d) $j=1$ , (e) $j=2$ and (f) total density; and of a circularly-asymmetric soliton for (g) $j=1$, (h) $j=2$ and (i) total density. In this figure and in Figs. 3, 4, 5 and 6 the energy is given in the plots of total density in the last column. The parameters are $c_{1}=c_{2}=0.25,\gamma=0.5$. The plotted quantities in all figures of this paper are dimensionless.
@@
Without losing generality we consider a nondipolar BEC with a small value of the nonlinearity parameters $c_{1}=c_{2}=0.25$ in Eqs. (II.1), so that a soliton can be formed for a small value of the SO-coupling strength $\gamma$, which facilitates both theoretical and numerical investigations. For a small $\gamma$ ($\gamma=0.5$) we find three types of degenerate solitons in a self-repulsive nondipolar BEC. First, there is a $(\mp 1,0)$-type multi-ring soliton for Rashba (upper sign) or Dresselhaus (lower sign) SO coupling. In Fig. 1 we display the contour plot of density of the central region of components (a) $j=1$, (b) $j=2$, and (c) total density of a $(\mp 1,0)$-type multi-ring soliton. The left (middle) column shows the density of component $j=1$ ($j=2$), whereas the right column illustrates the total density. The density is independent of whether we employ Dresselhaus SO coupling or Rashba SO coupling. However, the phase of this (+1,0)-type Dresselhaus SO-coupled soliton is distinct from a (1,0)(-1,0)-type Rashba SO-coupled soliton. In Figs. 2(a)-(b) we display the contour plot of the phase of wave function components $j=1,2$ of the $(+1,0)$-type Dresselhaus SO-coupled soliton of Figs. 1(a)-(c). In Fig. 2(a) there is a phase drop of $2\pi$ under a complete rotation, indicating an angular momentum projection of +1+1 (vortex of unit angular momentum) in component $j=1$ and 0 in component $j=2$. The corresponding phases for Rashba SO coupling are illustrated in Figs. 2(c)-(d) showing the phases of the wave-function components $j=1,2$ of the SO-coupled soliton. In this case in Fig. 2(c) we find a phase drop of $2\pi$ under a complete rotation around the origin, indicating an angular momentum projection of 1-1 (anti-vortex of unit angular momentum) in component $j=1$ and 0 in component $j=2$, viz. Fig. 2(d). In addition to the multi-ring solitons, for $\gamma=0.5$ we can also have stripe solitons with stripes in density, of a spatial period $\pi/\gamma$, of both the components as shown in Figs. 1(d)-(f) for components $j=1,2$ and total density, respectively. However, the total density of the two components is without any stripe pattern as shown in Fig. 1(f). Moreover, we have a circularly-asymmetric soliton as shown in Figs. 1(g)-(i), where we display the density of the two components and the total density, respectively. A close examination of the soliton components $j=1,2$ in Figs. 1(g)-(h), respectively, reveals the presence of asymmetrically located vortex-anti-vortex pairs, explicitly demonstrated through a contour plot of the phase of these wave-function components in Figs. 2(e)-(f). The component densities are zero (black) in Figs. 1(g)-(h) at the position of vortex and anti-vortex in Figs. 2(e)-(f), but the total density is not zero at these points, indicating that these are coreless. These three types of solitons are degenerate with the numerical energy ${\cal E}=-0.124$ close to its theoretical estimate ${\cal E}=\gamma^{2}/2=-0.125$ for $\gamma=0.5$. In this case no square-lattice soliton could be found. The scenario of soliton formation does not change for a self-attractive nondipolar BEC, where there are also these three types of soliton (result not presented in this paper). The degeneracy of the states depicted in Fig. 3, for small values of nonlinearities ($c_{1}=0.25,c_{2}=0.25$ and $d=0$), is a consequence of the degeneracy of the states in the linear problem in Sec. II.2, viz. Eqs. 20, (21)-(24). The degeneracy will disappear for a larger value of the nonlinearities.
@@
Figure 3: Contour plot of dimensionless density $n_{2D}(x,y)$ of (a)-(c) a square-lattice soliton, (d)-(f) a stripe soliton, and (g)-(i) a circularly-asymmetric soliton in a self-repulsive nondipolar SO-coupled BEC. The left [middle] column shows the density of component $j=1(n_{1})$ $j=2(n_{2})$], whereas the right column depicts the total density. The parameters of the model are $c_{1}=c_{2}=0.25$ and $\gamma=2$.
@@
As $\gamma$ is increased, a new scenario of soliton formation emerges. However, the types of solitons for $\gamma=1$ to $\gamma=3$ are quite similar and here we concentrate on $\gamma=2$. For $\gamma=2$, we find three types of degenerate solitons in a quasi-2D nondipolar self-repulsive SO-coupled BEC. The $(\mp 1,0)$-type multi-ring soliton, viz. Figs. 1(a)-(c), ceases to exist and gives rise to a new type of soliton: a square-lattice soliton with square-lattice modulation in density. This soliton is illustrated in Figs. 3(a)-(c) through a contour plot of the density of (a) component $j=1$, (b) component $j=2$, and (c) total density. The stripe soliton and the circularly-asymmetric soliton continue to exist. The resultant densities of the stripe soliton, with a spatial period of $\pi/\gamma$, are illustrated in Figs. 3(d)-(f). The densities of a circularly-asymmetric soliton are depicted in Figs. 3(g)-(i), which are very different from the same of the circularly-asymmetric soliton of Figs. 1(g)-(i) for $\gamma=0.5$. Also, there is no prominent vortex-anti-vortex structure in the wave functions of the two components as in the circularly-asymmetric soliton shown in Figs. 1(g)-(h) for $\gamma=0.5$, viz. Fig. 2(e)-(f). The stripe solitons of Figs. 1(d)-(f) and 3(d)-(f) are quite similar. However, in Figs. 3(a)-(f) the spatially periodic pattern has a small period consistent with $\gamma=2$. Although, there is a stripe pattern in density for both $\gamma=0.5$ and $\gamma=2$, the positions of maxima in component $j=1$ coincide with the minima in component $j=2$, thus resulting in a total density without modulation. For $\gamma=2$, the total density of the stripe soliton in Fig. 3(f) is identical to the same of the circularly-asymmetric soliton in Fig. 3(i), although the component densities are distinct. Hence, for a description of total density and energy, these two binary solitons behave like a unified whole, possessing the same total density and energy. We will see in the following that even in the presence of dipolar interaction the total densities of the stripe soliton and the circularly-asymmetric soliton continue to be equal. In this nondipolar case there is an $x\leftrightarrow y$ symmetry and there is a stripe and a circularly-asymmetric solution aligned along the $y$ axis with the same density and energy as the solitons in Figs. 3(d)-(f) and Figs. 3(g)-(i), respectively (not shown here). The solitons in Figs. 3(d)-(f) and Figs. 3(g)-(i), although are limited in the identical spatial region, they are distinct. The former has stripes (dark regions) along the $y$ axis; the area of the dark and the bright regions are the same. In the latter, there is no such dark region and the same number of atoms are distributed over the whole area. Consequently, the density in this case is half of the stripe soliton, compare the maximum densities 0.0024 and 0.0012 in the two cases. The numerical energy of the degenerate solitons illustrated in Fig. 3 is ${\cal E}=-1.999$ in perfect agreement with the theoretical estimate ${\cal E}=-\gamma^{2}/2$, which for $\gamma=2$ is ${\cal E}=-2$. For a self-attractive nondipolar SO-coupled BEC, we have the same three types of solitons (result not elaborated here).
@@ III.2 Quasi-2D SO-coupled binary self-repulsive and self-attractive dipolar BEC soliton
The scenario of soliton formation dramatically changes if we consider dipolar atoms with the polarization direction lying along the $y$ axis in the quasi-2D plane. This dipolar interaction is strongly anisotropic in the quasi-2D $x$-$y$ plane and it will break the circular symmetry of the system. With the introduction of the dipolar interaction, the SO-coupled self-repulsive dipolar system acquires additional attraction and the solitons become more compact and of smaller spatial dimension compared to the nondipolar case considered in Sec. III.1. For $\gamma=0.5$, the multi-ring, stripe and the circularly-asymmetric solitons continue to exist in this case, with much smaller size, as shown in Figs. 4(a)-(l), for $c_{1}=c_{2}=d=0.5$. Although, it is possible to have a quasi-2D dipolar soliton with the polarization direction in the quasi-2D plane malomed , there is no such soliton for these sets of parameters without an SO-coupling interaction. The dipolar nonlinearity $d=0.5$ is too weak to support such a soliton. The multi-ring soliton presented in Figs. 4(a)-(c) is of the ($\mp 1,0)$ type and has the phase structure similar to the same illustrated in Figs. 2(a)-(d) (not explicitly shown in this paper). The stripe soliton presented in Figs. 4(d)-(f) has a spatially periodic stripe along the $y$ axis. As in the nondipolar case, the two components of the spatially asymmetric soliton presented in Figs. 4(g)-(h) host a vortex-anti-vortex structure, which is explicit in the plot of the phase of these wave functions in Figs. 2(g)-(h). The total density and energy of the multi-ring soliton presented in Figs. 4(a)-(c) are the same as the circularly-asymmetric soliton illustrated in Figs. 4(g)-(i), although the component densities are distinct. Finally, we find a stripe soliton with stripe along the polarization $y$ axis shown in Figs. 4(j)-(l), which has practically the same density and also the energy as the stripe soliton with stripe along the $y$ axis, indicating an approximate $x\leftrightarrow y$ symmetry for a small $d=0.5$. All solitons of Fig. 4 are degenerate with energy ${\cal E}=-0.125$. We could not find any other type of soliton in this case.
Figure 4: Contour plot of dimensionless density $n_{2D}(x,y)$ of (a)-(c) a multi-ring soliton, (d)-(f) a stripe soliton with stripes along the $y$ axis, (g)-(i) a circularly-asymmetric soliton, and (j)-(l) a stripe soliton with stripes along the polarization $y$ axis. The left [middle] column shows the density of component $j=1(n_{1})$ $j=2(n_{2})$], whereas the right column depicts the total density. The parameters of the model are $c_{1}=c_{2}=d=0.5,$ and $\gamma=0.5$.
@@
Although, the principal interest of this paper is the formation of quasi-2D solitons in a self-repulsive SO-coupled dipolar BEC, it is also appropriate to investigate, if any, the changes in the structure of such solitons as the system changes from self-repulsive to self-attractive. To this end, in Fig. 6 we illustrate the densities of the quasi-2D solitons for a self-attractive SO-coupled dipolar BEC for $\gamma=2,c_{1}=c_{2}=-0.5,d=0.25.$ In this case, we reversed the sign of the strengths $c_{1}$ and $c_{2}$ of the contact interaction and reduced the same of the dipolar interaction $d$ from 0.5 to 0.25. A larger value of $d$ will imply more attraction, which will make the solitons to shrink to have a much smaller size. However, for this set of parameters ($c_{1}=c_{2}=-0.5,d=0.25$) there is no soliton in the absence of SO coupling; the dipolar nonlinearity ($d=0.25$) is too small to support a stable soliton. The circularly-asymmetric soliton of Figs. 5(d)-(f) is not possible in this case due to the added contact attraction. This soliton is eventually destroyed and turns to a stripe soliton with stripe along the $y$ axis. In this case, we find only two types of solitons: a stripe soliton with stripe along the $x$ axis with energy E=2.001E=-2.001 shown in Figs. 6(a)-(c) and a stripe soliton with stripe along the $y$ axis with energy E=2.008E=-2.008 depicted in Figs. 6(d)-(f). Due to an increased contact attraction the solitons of Fig. 6 are more compact and of much smaller size, when compared to the same in Fig. 5. In presence of stronger attraction, away from the linear limit, there is no degeneracy.
@@ IV Summary and Discussion
For a dipolar self-repulsive SO-coupled BEC, for a small SO coupling $\gamma=0.5,$ we find four types of degenerate solitons: a multi-ring soliton, a stripe soliton with stripe along the $x$ or $y$ axis, and a circularly-asymmetric soliton, as illustrated in Fig. 4. In the same system, for a large $\gamma(=2)$, we have three types of solitons: a stripe soliton with stripe along the $x$ or $y$ axis, and a circularly-asymmetric soliton aligned along the $y$ axis, viz. Fig. 5. The stripe solitons with the stripe along the $x$ axis and the polarization $y$ axis are degenerate with the circularly-asymmetric soliton emerging as an excited state. In a weakly self-attractive dipolar SO-coupled BEC, we find only two types of solitons. e.g. a stripe soliton with stripe along the $x$ axis or the $y$ axis. Of these two type the one with stripe along the $x$ axis is the ground state.

### v2

@@ III.1 Quasi-2D SO-coupled binary self-repulsive nondipolar BEC soliton
Figure 1: Contour plot of density $n_{j}$ of a quasi-2D nondipolar ($d=0$) self-repulsive multi-ring $(\mp 1,0)$-type binary Rashba or Dresselhaus SO-coupled BEC soliton for components (a) $j=1$ ($n_{1}$), (b) $j=2$ ($n_{2}$) and (c) total density ($n=n_{1}+n_{2}$); the same of a stripe soliton for components (d) $j=1$ , (e) $j=2$ and (f) total density; and of a circularly-asymmetric soliton for (g) $j=1$, (h) $j=2$ and (i) total density. In this figure and in Figs. 3, 4, 5 and 6 the energy is given in the plots of total density in the last column. The parameters are $c_{1}=c_{2}=0.25,\gamma=0.5$. The plotted quantities in all figures of this paper are dimensionless.
@@
Without losing generality we consider a nondipolar BEC with a small value of the nonlinearity parameters $c_{1}=c_{2}=0.25$ in Eqs. (II.1), so that a soliton can be formed for a small value of the SO-coupling strength $\gamma$, which facilitates both theoretical and numerical investigations. For a small $\gamma$ ($\gamma=0.5$) we find three types of degenerate solitons in a self-repulsive nondipolar BEC. First, there is a $(\mp 1,0)$-type multi-ring soliton for Rashba (upper sign) or Dresselhaus (lower sign) SO coupling. In Fig. 1 we display the contour plot of density of the central region of components (a) $j=1$, (b) $j=2$, and (c) total density of a $(\mp 1,0)$-type multi-ring soliton. The left (middle) column shows the density of component $j=1$ ($j=2$), whereas the right column illustrates the total density. The density is independent of whether we employ Dresselhaus SO coupling or Rashba SO coupling. However, the phase of this (+1,0)-type Dresselhaus SO-coupled soliton is distinct from a (1,0)(-1,0)-type Rashba SO-coupled soliton. In Figs. 2(a)-(b) we display the contour plot of the phase of wave function components $j=1,2$ of the $(+1,0)$-type Dresselhaus SO-coupled soliton of Figs. 1(a)-(c). In Fig. 2(a) there is a phase drop of $2\pi$ under a complete rotation, indicating an angular momentum projection of +1+1 (vortex of unit angular momentum) in component $j=1$ and 0 in component $j=2$. The corresponding phases for Rashba SO coupling are illustrated in Figs. 2(c)-(d) showing the phases of the wave-function components $j=1,2$ of the SO-coupled soliton. In this case in Fig. 2(c) we find a phase drop of $2\pi$ under a complete rotation around the origin, indicating an angular momentum projection of 1-1 (anti-vortex of unit angular momentum) in component $j=1$ and 0 in component $j=2$, viz. Fig. 2(d). In addition to the multi-ring solitons, for $\gamma=0.5$ we can also have stripe solitons with stripes in density, of a spatial period $\pi/\gamma$, of both the components as shown in Figs. 1(d)-(f) for components $j=1,2$ and total density, respectively. However, the total density of the two components is without any stripe pattern as shown in Fig. 1(f). Moreover, we have a circularly-asymmetric soliton as shown in Figs. 1(g)-(i), where we display the density of the two components and the total density, respectively. A close examination of the soliton components $j=1,2$ in Figs. 1(g)-(h), respectively, reveals the presence of asymmetrically located vortex-anti-vortex pairs, explicitly demonstrated through a contour plot of the phase of these wave-function components in Figs. 2(e)-(f). The component densities are zero (black) in Figs. 1(g)-(h) at the position of vortex and anti-vortex in Figs. 2(e)-(f), but the total density is not zero at these points, indicating that these are coreless. These three types of solitons are degenerate with the numerical energy ${\cal E}=-0.124$ close to its theoretical estimate ${\cal E}=\gamma^{2}/2=-0.125$ for $\gamma=0.5$. In this case no square-lattice soliton could be found. The scenario of soliton formation does not change for a self-attractive nondipolar BEC, where there are also these three types of soliton (result not presented in this paper).
@@
Figure 3: Contour plot of dimensionless density $n_{2D}(x,y)$ of (a)-(c) a square-lattice soliton, (d)-(f) a stripe soliton, and (g)-(i) a circularly-asymmetric soliton in a self-repulsive nondipolar SO-coupled BEC. The left [middle] column shows the density of component $j=1(n_{1})$ $j=2(n_{2})$], whereas the right column depicts the total density. The parameters of the model are $c_{1}=c_{2}=0.25,d=0$ and $\gamma=2$.
@@
As $\gamma$ is increased, a new scenario of soliton formation emerges. However, the types of solitons for $\gamma=1$ to $\gamma=3$ are quite similar and here we concentrate on $\gamma=2$. For $\gamma=2$, we find three types of degenerate solitons in a quasi-2D nondipolar self-repulsive SO-coupled BEC. The $(\mp 1,0)$-type multi-ring soliton, viz. Figs. 1(a)-(c), ceases to exist and gives rise to a new type of soliton: a square-lattice soliton with square-lattice modulation in density. This soliton is illustrated in Figs. 3(a)-(c) through a contour plot of the density of (a) component $j=1$, (b) component $j=2$, and (c) total density. The stripe soliton and the circularly-asymmetric soliton continue to exist. The resultant densities of the stripe soliton, with a spatial period of $\pi/\gamma$, are illustrated in Figs. 3(d)-(f). The densities of a circularly-asymmetric soliton are depicted in Figs. 3(g)-(i), which are very different from the same of the circularly-asymmetric soliton of Figs. 1(g)-(i) for $\gamma=0.5$. Also, there is no prominent vortex-anti-vortex structure in the wave functions of the two components as in the circularly-asymmetric soliton shown in Figs. 1(g)-(h) for $\gamma=0.5$, viz. Fig. 2(e)-(f). The stripe solitons of Figs. 1(d)-(f) and 3(d)-(f) are quite similar. However, in Figs. 3(a)-(f) the spatially periodic pattern has a small period consistent with $\gamma=2$. Although, there is a stripe pattern in density for both $\gamma=0.5$ and $\gamma=2$, the positions of maxima in component $j=1$ coincide with the minima in component $j=2$, thus resulting in a total density without modulation. For $\gamma=2$, the total density of the stripe soliton in Fig. 3(f) is identical to the same of the circularly-asymmetric soliton in Fig. 3(i), although the component densities are distinct. Hence, for a description of total density and energy, these two binary solitons behave like a unified whole, possessing the same total density and energy. We will see in the following that even in the presence of dipolar interaction the total densities of the stripe soliton and the circularly-asymmetric soliton continue to be equal. In this nondipolar case there is an $x\leftrightarrow y$ symmetry and there is a stripe and a circularly-asymmetric solution aligned along the $y$ axis with the same density and energy as the solitons in Figs. 3(d)-(f) and Figs. 3(g)-(i), respectively (not shown here). The solitons in Figs. 3(d)-(f) and Figs. 3(g)-(i), although are limited in the identical spatial region, they are distinct. The former has stripes (dark regions) along the $y$ axis; the area of the dark and the bright regions are the same. In the latter, there is no such dark region and the same number of atoms are distributed over the whole area. Consequently, the density in this case is half of the stripe soliton, compare the maximum densities 0.0024 and 0.0012 in the two cases. The numerical energy of the degenerate solitons illustrated in Fig. 3 is ${\cal E}=-1.999$ in perfect agreement with the theoretical estimate ${\cal E}=-\gamma^{2}/2$, which for $\gamma=2$ is ${\cal E}=-2$. The degeneracy of the states depicted in Fig. 3, for small values of nonlinearities ($c_{1}=0.25,c_{2}=0.25$ and $d=0$), is a consequence of the degeneracy of the states in the linear problem in Sec. II.2, viz. Eqs. 21, (22)-(25). The degeneracy will disappear for a larger value of the nonlinearities. For a self-attractive nondipolar SO-coupled BEC, we have the same three types of solitons (result not elaborated here).
@@
Figure 4: Contour plot of dimensionless density $n_{2D}(x,y)$ of (a)-(c) a multi-ring soliton, (d)-(f) a stripe soliton with stripes along the $y$ axis, (g)-(i) a circularly-asymmetric soliton, and (j)-(l) a stripe soliton with stripes along the $x$ axis. The left [middle] column shows the density of component $j=1(n_{1})$ $j=2(n_{2})$], whereas the right column depicts the total density. The parameters of the model are $c_{1}=c_{2}=d=0.5,$ and $\gamma=0.5$.
@@ III.2 Quasi-2D SO-coupled binary self-repulsive and self-attractive dipolar BEC soliton
The scenario of soliton formation dramatically changes if we consider dipolar atoms with the polarization direction lying along the $y$ axis in the quasi-2D plane. This dipolar interaction is strongly anisotropic in the quasi-2D $x$-$y$ plane and it will break the circular symmetry of the system. With the introduction of the dipolar interaction, the SO-coupled self-repulsive dipolar system acquires additional attraction and the solitons become more compact and of smaller spatial dimension compared to the nondipolar case considered in Sec. III.1. For $\gamma=0.5$, the multi-ring, stripe and the circularly-asymmetric solitons continue to exist in this case, with much smaller size, as shown in Figs. 4(a)-(l), for $c_{1}=c_{2}=d=0.5$. Although, it is possible to have a quasi-2D dipolar soliton with the polarization direction in the quasi-2D plane malomed , there is no such soliton for these sets of parameters without an SO-coupling interaction. The dipolar nonlinearity $d=0.5$ is too weak to support such a soliton. The multi-ring soliton presented in Figs. 4(a)-(c) is of the ($\mp 1,0)$ type and has the phase structure similar to the same illustrated in Figs. 2(a)-(d) (not explicitly shown in this paper). The stripe soliton presented in Figs. 4(d)-(f) has a spatially periodic stripe along the $y$ axis. As in the nondipolar case, the two components of the spatially asymmetric soliton presented in Figs. 4(g)-(h) host a vortex-anti-vortex structure, which is explicit in the plot of the phase of these wave functions in Figs. 2(g)-(h). The total density and energy of the multi-ring soliton presented in Figs. 4(a)-(c) are the same as the circularly-asymmetric soliton illustrated in Figs. 4(g)-(i), although the component densities are distinct. Finally, we find a stripe soliton with stripe along $x$ axis shown in Figs. 4(j)-(l), which has practically the same density and also the energy as the stripe soliton with stripe along the $y$ axis, indicating an approximate $x\leftrightarrow y$ symmetry for a small $d=0.5$. All solitons of Fig. 4 are degenerate with energy ${\cal E}=-0.125$. We could not find any other type of soliton in this case.
@@
Although, the principal interest of this paper is the formation of quasi-2D solitons in a self-repulsive SO-coupled dipolar BEC, it is also appropriate to investigate, if any, the changes in the structure of such solitons as the system changes from self-repulsive to self-attractive. To this end, in Fig. 6 we illustrate the densities of the quasi-2D solitons for a self-attractive SO-coupled dipolar BEC for $\gamma=2,c_{1}=c_{2}=-0.5,d=0.25.$ In this case, we reversed the sign of the strengths $c_{1}$ and $c_{2}$ of the contact interaction and reduced the same of the dipolar interaction $d$ from 0.5 to 0.25. A larger value of $d$ will imply more attraction, which will make the solitons to shrink to have a much smaller size. However, for this set of parameters ($c_{1}=c_{2}=-0.5,d=0.25$) there is no soliton in the absence of SO coupling; the dipolar nonlinearity ($d=0.25$) is too small to support a stable soliton. The circularly-asymmetric soliton of Figs. 5(d)-(f) is not possible in this case due to the added contact attraction. This soliton is eventually destroyed and turns to a stripe soliton with stripe along the $y$ axis. In this case, we find only two types of solitons: a stripe soliton with stripe along the $x$ axis with energy E=2.001E=-2.001 shown in Figs. 6(a)-(c) and a stripe soliton with stripe along the $y$ axis with energy E=2.008E=-2.008 depicted in Figs. 6(d)-(f). Due to an increased contact attraction, the solitons of Fig. 6 are more compact and of much smaller size, when compared to the same in Fig. 5. In the presence of a stronger attraction, away from the linear limit, there is no degeneracy of the solitons illustrated in Fig. 6.
@@ IV Summary and Discussion
For a dipolar self-repulsive SO-coupled BEC, for a small SO coupling $\gamma=0.5,$ we find four types of degenerate solitons: a multi-ring soliton, a stripe soliton with stripe along the $x$ or $y$ axis, and a circularly-asymmetric soliton, as illustrated in Fig. 4. In the same system, for a large $\gamma(=2)$, we have three types of solitons: a stripe soliton with stripe along the $x$ or $y$ axis, and a circularly-asymmetric soliton aligned along the $y$ axis, viz. Fig. 5. The stripe solitons with the stripe along the $x$ axis and the polarization $y$ axis are degenerate with the circularly-asymmetric soliton emerging as an excited state. In a weakly self-attractive dipolar SO-coupled BEC, we find only two types of solitons. e.g. a stripe soliton with stripe along the $x$ axis or the $y$ axis. Of these two type the one with stripe along the $y$ axis is the ground state, viz. Fig. 6.